1) Calculate the natural level of employment if the labor force is 150 million and the 

natural rate of unemployment is 5%.


1
Expert's answer
2022-01-17T12:29:47-0500

Natural rate of employment"=\frac {Employed}{Labor Force}"


"Natural rate of unemployment=\\frac {Unemployed}{Labor Force}\n\u200b"


"5 =\\frac {U}{150m}\\times100\n\u200b"


Where; U= Unemployed

"100U=5\times150m"

"U=\frac{(5\times 150m)}{100}=7.5m"

Employed = Labor force-Unemployed= 150-7.5= 142.5m

"Natural rate of employment=\frac {142.5m}{150m}\times 100= 95"%